# Advancing Multimodal Data Fusion in Pain Recognition: A Strategy Leveraging Statistical Correlation and Human-Centered Perspectives

### v1

@@ Abstract
This research tackles the challenge of integrating heterogeneous data for specific behavior recognition within the domain of Pain Recognition, presenting a novel methodology that harmonizes statistical correlations with a human-centered approach. By leveraging a diverse range of deep learning architectures, we highlight the adaptability and efficacy of our approach in improving model performance across various complex scenarios. The novelty of our methodology is the strategic incorporation of statistical relevance weights and the segmentation of modalities from a human-centric perspective, enhancing model precision and providing a explainable analysis of multimodal data. This study surpasses traditional modality fusion techniques by underscoring the role of data diversity and customized modality segmentation in enhancing pain behavior analysis. Introducing a framework that matches each modality with an suited classifier, based on the statistical significance, signals a move towards customized and accurate multimodal fusion strategies. Our contributions extend beyond the field of Pain Recognition by delivering new insights into modality fusion and human-centered computing applications, contributing towards explainable AI and bolstering patient-centric healthcare interventions. Thus, we bridge a significant void in the effective and interpretable fusion of multimodal data, establishing a novel standard for forthcoming inquiries in pain behavior recognition and allied fields. [^1] The code and model for this paper will be made publicly available upon acceptance.
@@ I Introduction
In the evolving domain of Human Computer Interaction, Affective Computing stands as a interdisciplinary field with pronounced relevance in enhancing human-computer interfaces and health surveillance[1]. It bridges emotional connections between humans and technology through the nuanced interpretation and simulation of human emotions, enhancing the capability of technology to identify and respond to emotional signals. Specifically, its role in understanding human pain behavior highlights its vast potential by recognizing pain as a complex emotional and not just a physical condition[2][3]. Studies reveal a significant Bayesian correlation between pain and anxiety, indicating a complex interaction between pain intensity, anxiety levels, and the potential indirect connection of defensive behaviors with anxiety[4].
@@
The complexity and multi-dimensionality of emotions and pain, pivotal in psychological and cognitive sciences, stem from the intricate interplay of factors within the nervous system, rather than isolated physiological or psychological causes[5]. This interplay, ingrained in our neurophysiological structure, illustrates that emotions are the product of diverse factors, underscored by neural processes[6][7].
The analysis of pain’s complexity gains unparalleled insights through multimodal data analysis. Advancements in human-centered computing have been propelled by multimodal data fusion, significantly enhancing emotional recognition and interactivity. Despite these advancements, challenges remain, particularly in dealing with data heterogeneity and alignment issues across different modalities[8]. Deep learning offers improvements in data processing but faces interpretability and high-dimensional data integration challenges[9]. Traditional approaches sometimes neglect the complexity and ethical considerations inherent in human-centered computing and pain evaluation. Emphasizing technological efficiency, applicability, and ethics, human-centered computing necessitates integrating human factors like physiology, behavior, emotions, and culture into model development, thereby ensuring AI models’ interpretability and ethicality[10][2].
This investigation explores the effectiveness of statistical methods in multimodal data analysis, focusing on predicting protective behaviors with significant implications for pain recognition. Despite some successes of conventional machine learning and deep learning techniques in analyzing pain behaviors, they are limited in precision and interpretability amidst the complexities of highly heterogeneous multimodal data[11]. This research aims to improve the accuracy and utility of pain analysis by optimizing modality fusion through statistical methods and incorporating a data-driven, human-centric approach. The statistical approach helps in dissecting and understanding the intricate relationships among different modalities, thus facilitating a more accurate analysis of pain behaviors, identifying essential modalities for the enhancement of predictive models, and reducing redundancies[12]. Through statistical inference and hypothesis testing, this study validates the contributions of individual modalities, allowing for dynamic weight adjustments to improve model adaptability and personalization. This not only augments the precision, efficiency, and interpretability of the models but also addresses ethical considerations, steering towards more effective analyses of pain behaviors and the development of personalized, precision pain management strategies.
@@ II Related Work
Previous studies suggested that traditional machine learning and deep learning approaches could present an alternative to manual pain scoring, such as using a conventional neural network, based on Google’s InceptionV3 model, to predict a binary label of “pain” and “no pain” from mouse face images [13]. In addition to this, Surface ElectroMyoGraphy (SEMG) recordings have been identified as a valuable tool for assessing chronic low back pain [14]. Building upon this foundation of research, there emerges a novel perspective within pain recogniction, wherein the integration of multimodal data can use the strengths of various sources of data to enhance the accuracy and precision of predicting pain-related behavior. However, to begin, it is essential to establish and evaluate pain as a human emotion. Past research concluded that there are overlapping regions of the brain in the Central Nervous System (CNS) between pain and other emotions, encompassing areas such as the amygdala, the thalamus, and the Anterior Cingulate Cortex (ACC) [15].
@@
This complexity necessitates an integrated, multimodal analytical approach that goes beyond singular data sources. Consequently, state-of-the-art deep learning algorithms, including CNN-LSTMs, have been leveraged to significantly improve the accuracy and speed of pain detection [16][17]. However, the challenge of effectively integrating and processing this heterogeneous data persists, and overcoming this challenge is essential for advancing multimodal data fusion and maximizing the efficacy of analytical methods for pain identification [11].
@@
Reflecting on this necessity, historical research reinforces the value of integrating multiple data types, such as body movements and muscle activity, to enhance pain detection accuracy [18]. These findings highlight the crucial interconnectivity of different modalities, underpinning their essential role in pain analysis and supporting the advancement of multimodal methodologies in this study.
@@
Subsequent explorations have further elucidated this concept, with studies illustrating the efficacy of combining muscle and motion signals for the accurate identification of protective behaviors—a key aspect of pain behavior analysis. Notably, central (model-level) fusion approaches have been shown to outperform both feature and decision-level fusion methods in the context of Protective Behavior Detection (PBD), showcasing their superior capability in harnessing the potential of multimodal data[19][20]. This advancement signals a critical insight into the differential impact of various fusion layers on the process of pain recognition, highlighting the imperative of selecting and implementing the most effective fusion strategies to achieve a more integrative and holistic understanding of pain behaviors through the synthesis of multimodal information.
@@
Innovative model architectures have been proposed to improve the precision and applicability of pain analysis studies. The P-STEMR framework is one such innovation, which capitalizes on human activity recognition datasets for the categorization of pain levels, illuminating the capabilities of supervised learning in scenarios where labeled datasets are scarce[21]. Concurrently, the advent of a cutting-edge hierarchical HAR-PBD architecture signifies a step forward in enabling real-time pain assessment. This architecture synergizes Human Activity Recognition (HAR) with Protective Behavior Detection (PBD) through the integration of graph convolution and Long Short-Term Memory (LSTM) networks. Furthermore, this model innovatively addresses the prevalent issue of data classification imbalance by incorporating the Class-Balanced Focal Classification Cross-Entropy (CFCC) loss mechanism [22], demonstrating a methodical approach to improve the reliability and effectiveness of pain monitoring systems.
@@
The challenges in multimodal data fusion lie in the seamless integration of heterogeneous data from varied sources for effective pain state recognition, where increased dimensionality muddles model interpretability. Although previous studies have leveraged machine learning and deep learning to automate pain scoring to some extent, they fall short when dealing with complex, multi-source pain behaviors. Our innovative approach underscores a multimodal data fusion strategy anchored in statistical relevance, enriched by a human-centric perspective to bolster model efficacy. This methodology not only acknowledges the diversity of data but ingeniously converts the correlations between various modal data and outcomes into weighted contributions for model decision-making, thereby ensuring a more precise and holistic detection of pain behaviours.
@@ III Methodology
The pivotal inquiry of this research concentrates on devising methodologies for the effective fusion and processing of multimodal data derived from varied sources and types. This encompasses leveraging statistical tools for the fusion of modal features, conducting hypothesis tests on target variables, and performing correlation analyses. Moreover, this study is dedicated to exploring the evaluation of modal significance and the distribution of dynamic weights predicated on statistical findings, in addition to determining methodologies for optimizing the combination of features to enhance the efficacy of models predicting pain.
@@ III-A Dataset Introduction
In this research, we harnessed the EmoPain dataset, an instrumental resource for investigating the correlation between body movements and levels of pain intensity[23]. This dataset is bifurcated into training and validation sets, incorporating data from 10 chronic pain sufferers and 6 healthy controls in the former, and 4 chronic pain individuals alongside 3 healthy controls in the latter. In Table III-A and Figure 2, highlighted within the dataset are the X, Y, and Z coordinates of body joints, segmented across columns 1-22, 23-44, and 45-66, respectively. The crux of our analysis is centered on vector 73, which gauges protective behavior, identifying non-protective actions with 0 and protective behaviors with 1. This intricate relationship between protective behaviors and pain, intermediated by emotional states, situates our study at the convergence of behavioral analysis, pain evaluation, and emotional computation, suggesting significant repercussions for the domain of human-centered computing.
@@ III-B Dataset Analysis
The histograms and Q-Q plots in Figure LABEL:fig:traindis and Figure LABEL:fig:validdis for the training and validation data provide crucial insights into the normality of the dataset’s distribution. From the histograms, both datasets display a pronounced peak, but it doesn’t align with the theoretical normal distribution curve, suggesting a discrepancy from normality. This is further evidenced by the asymmetry and apparent deviations from the central peak.
@@
Given the distribution characteristics of our dataset, as outlined in Table I, we identify the limitations and suitability of various statistical methods for our analysis. The assumption of normality renders ANOVA and Pearson correlation ineffective for our purposes, while Kendall’s rank correlation, despite its robustness, is impractical due to computational demands and is more suited for smaller datasets. Consequently, we opt for Spearman’s rank correlation as the most fitting choice, owing to its efficiency with large datasets and applicability to non-linear relationships, ensuring a more accurate and relevant analysis of our multimodal data. This decision strategically aligns with our analytical needs, emphasizing Spearman’s correlation as the most effective tool for deciphering the complex relationships within our heterogeneous dataset.
@@ III-C5 Comparative Analysis & Model Performance Evaluation
By juxtaposing the performance of the singular modality (benchmark model) with both the bifurcated and quadrifurcated modality approaches, as well as the average weighted voting benchmark, the experiment aspires to illuminate whether a granular, human-centered feature segmentation supplemented by statistically weighted decision-level fusion markedly optimizes model accuracy in pain evaluation. The transition from a singular, homogenously weighted modality to a nuanced, statistically or evenly weighted integration of multiple modalities endeavors to elucidate the symbiotic relationship between data-driven and human-centered segmentation, and their collective prowess in enhancing pain recognition models. This comprehensive exploration aims to crystallize the efficacy of each strategic approach and underscore their cumulative contribution towards refining model performance in the nuanced domain of pain analysis.
@@ IV Results and Evaluation
Our findings underscore the pivotal role of modality segmentation and decision-level fusion, informed by statistical insights, in augmenting the accuracy of protective behavior detection. They also highlight the broad applicability of statistically driven weighting strategies within various complex neural frameworks. Through an in-depth analysis of performance shifts across different configurations, we reveal the critical role of integrating sophisticated data processing techniques, human-cenered perspectives, and statistical correlations. This holistic strategy markedly enhances the precision and effectiveness of protective behavior recognition in complex physiological datasets, representing a significant advance toward nuanced patient-centered healthcare solutions. Given the imbalanced nature of our dataset, we prioritize precision, recall, and F1 score over accuracy to more accurately reflect model performance across labels. We employed four distinct models— LSTM, CNN, CNN-Attention, and CNN with Multi-head Self-Attention—to diversify our evaluation and mitigate single-model reliance risks. It is important to note that a comparative performance analysis among these models is not within this paper’s scope. This methodology underscores our dedication to a thorough evaluation, aiming to deepen the understanding of multi-modal data fusion’s impact on model efficacy amidst dataset imbalances.
The data presented in Table II offer a detailed examination of performance variances across models utilizing different modality integrations. A notable observation is that models based on a single modality can achieve high accuracy, reaching up to 0.93, but this is often at the sacrifice of precision, recall, and F1-score, which tend to remain around 0.55. This discrepancy highlights the potential trade-off between optimizing for accuracy and maintaining balanced performance across all metrics. Investigating models that incorporate dual modalities further reveals differential effects on performance metrics. For instance, transitioning to a dual-modality configuration results in a slight reduction in accuracy for the LSTM model, contrasting with improvements observed in other architectures. This variation underscores the potential limitations of relying solely on a single modality, which, while yielding high accuracy, may not effectively capture broader evaluative measures. In contrast, the adoption of an additional modality leads to significant enhancements in precision, recall, and F1-scores, with increases ranging from 12% to 62%. This strongly supports the premise that integrating multiple modalities substantially bolsters model robustness. Segregating input features into distinct spatial and sEMG modalities, rather than amalgamating all 70 features, proves strategically beneficial. This approach amplifies the model’s ability to accurately identify and assess protective behaviors, demonstrating improved precision and sensitivity.
@@
Adopting a human-centered approach to segment the dataset into four modalities demonstrates potential enhancements in model performance. Particularly, models employing CNN-Attention mechanisms exhibit slight but positive differences, suggesting an improved capability in capturing protective behavior features. This indicates that modality segmentation, guided by human-centered principles, can amplify model effectiveness. Further analysis reveals that models employing statistical weighting generally outperform those using mean weighting, except in the case of LSTM with four modalities employing average weighting—a scenario that mirrors the trade-offs observed in single-modality LSTM models.
This investigation underscores the advantage of multimodal strategies, where strategic feature grouping and statistical decision-making markedly elevate model efficacy. Our findings particularly highlight the utility of attention mechanisms, like CNN-Attention and CNN with Multi-head Self-Attention, in a four-modality framework, reinforcing the benefits of human-centered modality segmentation. Overall, the transition to a four-modality model, grounded in human-centered design and statistical weighting, is validated as superior to single-modality approaches, bolstering the case for sophisticated multimodal fusion in enhancing protective behavior recognition.
The examination distinctly emphasizes the superiority of statistical weighting over average weighting in multimodal configurations, enhancing model performance, with an exception observed in the LSTM model utilizing average weighting, suggesting a nuanced balance between accuracy and other metrics. Despite this, the evidence strongly supports the multimodal approach’s advantages over singular modality frameworks, reinforcing that strategic segmentation of modalities, when combined with statistical weighting, markedly advances protective behavior detection models.

### v2

# Advancing Pain Recognition through Statistical Correlation-Driven Multimodal Fusion*

@@ Abstract
This research presents a novel multimodal data fusion methodology for pain behavior recognition, integrating statistical correlation analysis with human-centered insights. Our approach introduces two key innovations: 1) integrating data-driven statistical relevance weights into the fusion strategy to effectively utilize complementary information from heterogeneous modalities, and 2) incorporating human-centric movement characteristics into multimodal representation learning for detailed modeling of pain behaviors. Validated across various deep learning architectures, our method demonstrates superior performance and broad applicability. We propose a customizable framework that aligns each modality with a suitable classifier based on statistical significance, advancing personalized and effective multimodal fusion. Furthermore, our methodology provides explainable analysis of multimodal data, contributing to interpretable and explainable AI in healthcare. By highlighting the importance of data diversity and modality-specific representations, we enhance traditional fusion techniques and set new standards for recognizing complex pain behaviors. Our findings have significant implications for promoting patient-centered healthcare interventions and supporting explainable clinical decision-making.
@@ I Introduction
Affective Computing, an interdisciplinary field within Human-Computer Interaction, holds immense potential for enhancing human-computer interfaces and health surveillance by recognizing and responding to human emotions[1]. Its application in understanding human pain behavior is particularly promising, as it acknowledges pain as a complex emotional state rather than merely a physical condition [2][3]. As illustrated in Figure 1, pain is closely intertwined with anxiety, which can lead to protective behaviors such as guarding[4]. However, accurate pain recognition remains challenging due to the multi-dimensionality and subjectivity of pain experiences, which stem from the intricate interplay of physiological, psychological, and social factors[5][6][7].
@@
Multimodal data fusion has emerged as a powerful approach in human-centered computing to address these challenges by integrating information from diverse sources such as physiological signals, behavioral cues, and self-reports[8]. Nevertheless, existing multimodal pain recognition methods often struggle with data heterogeneity, alignment issues, and limited interpretability[9]. Moreover, conventional machine learning and deep learning techniques often fail to capture the complexities of pain experiences and lack the adaptability to personalize pain recognition[10].
To overcome these limitations, this research introduces a novel approach to multimodal pain recognition that synergistically integrates statistical methods with a human-centric perspective. Our key innovations lie in: 1) employing statistical inference and hypothesis testing to explore the relationships between different modalities and the target variable (pain states), identifying the most informative features for pain recognition; and 2) incorporating human-centered insights into the representation and modeling of pain experiences to ensure the interpretability and ethical alignment of our models. By dynamically adjusting the contributions of different modalities through adaptive weighting, our data-driven, personalized approach enhances the precision, efficiency, and adaptability of pain recognition. Ultimately, this research contributes to the advancement of pain recognition, affective computing, AI-assisted healthcare, and the development of empathetic and socially responsible AI systems, paving the way for more effective and ethically grounded pain management strategies.
@@ II Related Work
Previous studies have suggested that traditional machine learning and deep learning approaches can serve as alternatives to manual pain scoring. For instance, using a conventional neural network based on Google’s InceptionV3 model, researchers predicted binary labels of “pain” and “no pain” from mouse facial images [11]. Additionally, Surface Electromyography (SEMG) recordings have proven valuable in assessing chronic low back pain [12]. Building on this foundational research, a new perspective in pain recognition emphasizes the integration of multimodal data to leverage the strengths of various sources, thereby enhancing the accuracy and precision of predicting pain-related behaviors. However, it is crucial first to establish and evaluate pain as a human emotion. Previous research has identified overlapping brain regions within the Central Nervous System (CNS) that are involved in both pain and other emotions, including the amygdala, thalamus, and Anterior Cingulate Cortex (ACC) [13].
@@
This complexity necessitates an integrated, multimodal analytical approach that goes beyond singular data sources. Consequently, state-of-the-art deep learning algorithms, including CNN-LSTMs, have been leveraged to significantly improve the accuracy and speed of pain recognition [14][15]. However, the challenge of effectively integrating and processing this heterogeneous data persists, and overcoming this challenge is essential for advancing multimodal data fusion and maximizing the efficacy of analytical methods for pain recognition [10].
@@
Reflecting on this necessity, historical research reinforces the value of integrating multiple data types, such as body movements and muscle activity, to enhance pain recognition accuracy [16]. These findings highlight the crucial interconnectivity of different modalities, underpinning their essential role in pain recognition and supporting the advancement of multimodal methodologies in this study.
@@
Subsequent explorations have further elucidated this concept, with studies illustrating the efficacy of combining muscle and motion signals for the accurate identification of protective behaviors—a key aspect of pain recognition. Notably, central (model-level) fusion approaches have been shown to outperform both feature and decision-level fusion methods in the context of Protective Behavior Detection (PBD), showcasing their superior capability in harnessing the potential of multimodal data[17][18]. This advancement signals a critical insight into the differential impact of various fusion layers on the process of pain recognition, highlighting the imperative of selecting and implementing the most effective fusion strategies to achieve a more integrative and holistic understanding of pain behaviors through the synthesis of multimodal information.
@@
Innovative model architectures have been developed to enhance the accuracy and applicability of pain recognition studies. The P-STEMR framework is a notable example, utilizing human activity recognition datasets to classify pain levels, showcasing the utility of supervised learning in situations with limited labeled data sources [19]. Additionally, the introduction of an advanced hierarchical HAR-PBD architecture represents significant progress in real-time pain recognition. This architecture combines Human Activity Recognition (HAR) with Protective Behavior Detection (PBD) using graph convolution and Long Short-Term Memory (LSTM) networks. Moreover, this model addresses the common challenge of data classification imbalance by implementing the Class-Balanced Focal Classification Cross-Entropy (CFCC) loss function [20], thereby enhancing the reliability and effectiveness of pain monitoring systems.
@@
The challenges in multimodal data fusion involve integrating heterogeneous data from diverse sources for accurate pain state recognition, where increased dimensionality can reduce model interpretability. While previous research has employed machine learning and deep learning for automated pain recognition, these methods often struggle with complex, multi-source pain behaviors. Our novel approach emphasizes a multimodal data fusion strategy grounded in statistical relevance, complemented by a human-centric perspective to enhance model effectiveness. This methodology not only accounts for data diversity but also transforms the relationships between different modalities and outcomes into weighted contributions for model decision-making, resulting in more precise and comprehensive detection of pain behaviors.
@@ III Methodology
The central focus of this research is on developing methodologies for the effective fusion and processing of multimodal data from diverse sources and types. This involves utilizing statistical techniques to integrate modal features, conducting hypothesis tests on target variables, and performing correlation analyses. Additionally, the study aims to evaluate the significance of different modalities and determine dynamic weight distribution based on statistical insights. It also seeks to optimize the combination of features to improve the accuracy and effectiveness of models in predicting pain.
@@ III-A Dataset Introduction
In this research, we utilized the EmoPain dataset, a key resource for exploring the relationship between body movements and pain intensity levels [21]. The dataset is divided into training and validation sets, with data from 10 chronic pain sufferers and 6 healthy controls in the training set, and 4 chronic pain individuals and 3 healthy controls in the validation set. As detailed in Table III-A and Figure 2, the dataset includes X, Y, and Z coordinates of body joints, categorized in columns 1-22, 23-44, and 45-66, respectively. The core of our analysis focuses on vector 73, which measures protective behavior, distinguishing non-protective actions (coded as 0) from protective behaviors (coded as 1). This complex interplay between protective behaviors and pain, mediated by emotional states, positions our study at the intersection of behavioral analysis, pain recognition, and emotional computation.
@@ III-B Dataset Analysis
The histograms and Q-Q plots in Figure 3 for the training and validation data provide crucial insights into the normality of the dataset’s distribution. From the histograms, both datasets display a pronounced peak, but it doesn’t align with the theoretical normal distribution curve, suggesting a discrepancy from normality. This is further evidenced by the asymmetry and apparent deviations from the central peak.
@@
Given the distribution characteristics of our dataset, as outlined in Table I, we identify the limitations and suitability of various statistical methods for our analysis. The assumption of normality renders ANOVA and Pearson correlation ineffective for our purposes, while Kendall’s rank correlation, despite its robustness, is impractical due to computational demands and is more suited for smaller datasets. Consequently, we opt for Spearman’s rank correlation as the most fitting choice, owing to its efficiency with large datasets and applicability to non-linear relationships, ensuring a more accurate and relevant analysis of our multimodal data. This decision strategically aligns with our analytical needs, highlighting Spearman’s correlation as the optimal tool for exploring the complex relationships within our heterogeneous dataset.
@@ III-C5 Comparative Analysis & Model Performance Evaluation
By comparing the performance of the singular modality (benchmark model) with both the bifurcated and quadrifurcated modality approaches, as well as the average weighted voting benchmark, the experiment aspires to illuminate whether a granular, human-centered feature segmentation supplemented by statistically weighted decision-level fusion markedly optimizes model accuracy in pain recognition. The transition from a singular, homogenously weighted modality to a nuanced, statistically or evenly weighted integration of multiple modalities endeavors to elucidate the symbiotic relationship between data-driven and human-centered segmentation, and their collective prowess in enhancing pain recognition models. This comprehensive exploration aims to crystallize the efficacy of each strategic approach and underscore their cumulative contribution towards refining model performance in the nuanced domain of pain recognition.s
@@ IV Results and Evaluation
Our findings emphasize the pivotal role of modality segmentation and decision-level fusion, informed by statistical insights, in augmenting the accuracy of protective behavior detection. They also highlight the broad applicability of statistically driven weighting strategies within various complex neural frameworks. Through an in-depth analysis of performance shifts across different configurations, we reveal the critical role of integrating sophisticated data processing techniques, human-centered perspectives, and statistical correlations. This holistic strategy markedly enhances the precision and effectiveness of protective behavior recognition in complex physiological datasets, representing a significant advance toward nuanced patient-centered healthcare solutions. Given the imbalanced nature of our dataset, we prioritize precision, recall, and F1 score over accuracy to more accurately reflect model performance across labels. We employed four distinct models— LSTM, CNN, CNN-Attention, and CNN with Multi-head Self-Attention—to diversify our evaluation and mitigate single-model reliance risks. It is important to note that a comparative performance analysis among these models is not within this paper’s scope. This methodology underscores our dedication to a thorough evaluation, aiming to deepen the understanding of multi-modal data fusion’s impact on model efficacy amidst dataset imbalances.
The data presented in Table II provide a detailed analysis of performance differences across models using various modality integrations. A key observation is that models based on a single modality can achieve high accuracy, up to 0.93, but often at the expense of other metrics like precision, recall, and F1-score, which hover around 0.55. This indicates a potential trade-off between maximizing accuracy and ensuring balanced performance across all metrics. Examining models with dual modalities reveals different impacts on performance metrics. For example, while the LSTM model experiences a slight drop in accuracy with dual-modality integration, other architectures show improvements. This variation highlights the limitations of relying solely on a single modality, which, despite high accuracy, may not fully capture a broader range of evaluative measures. Conversely, incorporating an additional modality significantly enhances precision, recall, and F1-scores, with increases ranging from 12% to 62%. This strongly supports the notion that integrating multiple modalities enhances model robustness. Specifically, segregating input features into distinct spatial and sEMG modalities, rather than combining all 70 features, proves to be a strategic advantage. This approach improves the model’s ability to identify and evaluate protective behaviors, resulting in better precision and sensitivity.
@@
Adopting a human-centered approach to segment the dataset into four modalities demonstrates potential enhancements in model performance. Particularly, models employing CNN-Attention mechanisms exhibit slight but positive differences, suggesting an improved capability in capturing pain behavior features. This indicates that modality segmentation, guided by human-centered principles, can amplify model effectiveness. Further analysis reveals that models employing statistical weighting generally outperform those using mean weighting, except in the case of LSTM with four modalities employing average weighting—a scenario that mirrors the trade-offs observed in single-modality LSTM models.
This investigation emphasizes the advantage of multimodal strategies, where strategic feature grouping and statistical decision-making markedly elevate model efficacy. Our findings particularly highlight the utility of attention mechanisms, like CNN-Attention and CNN with Multi-head Self-Attention, in a four-modality framework, reinforcing the benefits of human-centered modality segmentation. Overall, the transition to a four-modality model, grounded in human-centered design and statistical weighting, is validated as superior to single-modality approaches, bolstering the case for sophisticated multimodal fusion in enhancing protective behavior recognition.
The examination distinctly emphasizes the superiority of statistical weighting over average weighting in multimodal configurations, enhancing model performance. However, an exception was observed in the LSTM model utilizing average weighting, suggesting a nuanced balance between accuracy and other metrics. Despite this, the evidence strongly supports the advantages of the multimodal approach over singular modality frameworks. This finding reinforces the idea that strategic segmentation of modalities, when combined with statistical weighting, markedly advances protective behavior detection models.